INAF queue observing, 2015 Jan 10-11

Observer: E. Sani, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Michelle Edwards
Instrument: MODS, LUCI

Summary

At the beginning of the night the dome stays closed due to humidity, the temperature reaches the dew point. Humidity 100%, dewpoint -2.4, external temperature -2.6. No wind.

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI Calibrations

Details

01:15 We start LUCI flats and arcs
  • executeLUCIScript.sh A2_LS_Ks.flat luci.0001-0010 flats counts are around 30000-35000,
  • we run some test manually on the Read Out Manager Gui and in the end repeat the execution with HALO2 lamp and 10s of exposure time
    • executeLUCIScript.sh A2_LS_Ks.flat luci.0036-0045

  • luci.0047-0048 are not good

  • executeLUCIScript.sh Target_11D.flat luci.0049-0058 Good also for Target_11F
  • executeLUCIScript.sh Target_11E.flat luci.0059-0068

-- %USERSIG{EleonoraSani - 2015-01-12}%

Comments

 
Topic revision: r6 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback